# Hierarchical Open-Vocabulary 3D Scene Graphs for Language-Grounded Robot Navigation

### v1

@@ Abstract
Recent open-vocabulary robot mapping methods enrich dense geometric maps with pre-trained visual-language features. While these maps allow for the prediction of point-wise saliency maps when queried for a certain language concept, large-scale environments and abstract queries beyond the object level still pose a considerable hurdle, ultimately limiting language-grounded robotic navigation. In this work, we present HOV-SG, a hierarchical open-vocabulary 3D scene graph mapping approach for language-grounded robot navigation. Leveraging open-vocabulary vision foundation models, we first obtain state-of-the-art open-vocabulary segment-level maps in 3D and subsequently construct a 3D scene graph hierarchy consisting of floor, room, and object concepts, each enriched with open-vocabulary features. Our approach is able to represent multi-story buildings and allows robotic traversal of those using a cross-floor Voronoi graph. HOV-SG is evaluated on three distinct datasets and surpasses previous baselines in open-vocabulary semantic accuracy on the object, room, and floor level while producing a 75% reduction in representation size compared to dense open-vocabulary maps. In order to prove the efficacy and generalization capabilities of HOV-SG, we showcase successful long-horizon language-conditioned robot navigation within real-world multi-storage environments. We provide code and trial video data at: https://hovsg.github.io.
@@ I Introduction
Humans acquire conceptual knowledge through multi-modal experiences. These experiences are paramount to object recognition and language as well as reasoning and planning [1, 2]. Cognitive maps store this information based on sensor fusion, fragmentation, and hierarchical structure. This is central to the human ability to navigate the physical world [3, 4, 5]. Recently, language proved to be an effective link between intelligent systems and humans and can enable robot autonomy in complex human-centered environments [6, 7, 8, 9, 10, 11, 12, 13].
Classical methods for robot navigation build dense spatial maps of high accuracy using approaches to simultaneous localization and mapping (SLAM) [14, 15]. Those give rise to fine-grained navigation and manipulation based on geometric goal specifications. Recent advances have combined dense maps with pre-trained zero-shot vision-language models, which facilitates open-vocabulary indexing of observed environments [6, 7, 8, 16, 10, 17, 9]. While these approaches marry the area of classical robotics with modern open-vocabulary semantics, representing larger scenes while abstracting still poses a considerable hurdle. Scalable scene representations generated from real-world perception inputs should generally fulfill the following requirements: 1) Object-centricity and abstraction in terms of hierarchies, 2) efficiency regarding storage capacity as well as efficacy in terms of actionability, 3) true open-vocabulary indexing and simple querying.
@@
In this work, we present Hierarchical Open-Vocabulary 3D Scene Graphs, short HOV-SG. Our approach abstracts from dense panoptic maps and allows indexing of three distinct concepts, namely floors, rooms as well as objects. We utilize open-vocabulary vision-language models [20, 21] across all concepts in order to construct 3D scene graph hierarchies that span multi-story environments while maintaining a small memory footprint. Given its concept-centric nature, the HOV-SG representation is promptable using LLMs. Different from previous work, our approach relates to different conceptual levels by first decomposing abstract queries such as “towel in the bathroom on the upper floor” and scoring the obtained tokens against the different levels of the hierarchy. We complement this with a navigational Voronoi graph that covers multiple floors including stairs, which allows actionable grounding of decomposed queries in the environment. This enables object retrieval and long-horizon robotic navigation in large-scale indoor environments from abstract queries as shown in Fig. 1.
As part of this work, we present state-of-the-art results in open-set 3D semantic segmentation, object retrieval from long queries as well as robust semantic room classification from perception inputs. In addition, we propose a novel open-vocabulary semantic similarity metric termed AUC. It measures the area under the top-k accuracy curve and thus scales to very large numbers of object categories and to differently-sized label sets used for evaluation. In summary, we make the following main contributions:
We introduce a pipeline for constructing hierarchical, open-vocabulary 3D scene graphs from multi-story environments that enables LLM-based prompting, planning as well as robotic navigation.
We extensively evaluate our approach across three diverse datasets as well as a real-world environment. We achieve state-of-the-art results in 3D open-set semantic segmentation and demonstrate successful robotic navigation from abstract language queries.
We introduce a novel evaluation metric for measuring open-vocabulary semantics termed AUC and publish code at https://hovsg.github.io.
@@ II-A Semantic 3D Mapping
3D scene graphs have emerged as an effective, object-centric representation of large-scale indoor [32, 18, 19] and outdoor scenes [15]. By representing objects or spatial concepts as nodes and their relations as edges, 3D scene graphs allow to efficiently represent larger scenes [32, 15, 18]. Both edges and nodes can hold geometric and semantic attributes, which are often inferred from certain off-the-shelf networks [33]. Decomposing scenes into objects and their relations enables higher-level reasoning for robotic navigation and manipulation. This is particularly useful in the realm of reasoning, planning, and navigation given the object-centric nature of these tasks [18, 34, 35]. Often combined with odometry estimates from simultaneous localization and mapping (SLAM) [36, 15, 37], 3D scene graphs also allow a tight coupling between semantics and highly accurate mapping approaches utilizing e.g. meshes to represent environments [18].
@@
Early works have shown how to encode hierarchies via abstraction in both the spatial and the semantic domain using offline approaches [32, 19]. Successive works investigated learning-based scene graph construction [38, 33] as well as dynamic indoor scenes [18]. Several approaches such as SceneGraphFusion [33] and S-Graphs [36] also investigate the real-time capabilities of their proposed approaches. Most recently, ConceptGraphs [11] was the first to show how to combine 3D scene graphs with open-vocabulary vision-language features. In addition, the authors show how to query the graph using LLMs and demonstrate interesting application scenarios.
@@
Conceptually, our work is most similar to ConceptGraphs and Hydra. While ConceptGraphs is only evaluated on small scenes and mostly validated by human evaluators in terms of semantic accuracy of nodes etc. Different from ConceptGraphs and similar to Hydra [18], which does not operate on open-vocabulary features, our approach demonstrates how to efficiently represent actionable, hierarchical 3D scene graphs that are attributed with open-vocabulary features. These scene graph hierarchies span entire multi-story environments and allow robotic navigation and search from abstract queries. In addition to this, we outperform ConceptGraphs in terms of 3D semantic segmentation accuracy and make contributions regarding the open-vocabulary semantic classification of room types. Moreover, the proposed open-vocabulary graph representation of the environment is considerably more memory efficient by consuming an average of 75% less storage compared to previous dense open-vocabulary representations while preserving extensive semantic information.
@@ III Technical Approach
This work aims to develop a concise and efficient visual-language graph representation for large-scale multi-floor indoor environments given RGB-D observations and odometry. The graph should facilitate the indexing of semantic concepts through natural language queries and enable robotic navigation in multi-floor environments. We address this by proposing Hierarchical Open-Vocabulary Scene Graphs, in short HOV-SG. In the following, we describe (i) the construction of a 3D segment-level open-vocabulary map (Sec. III-A), (ii) the generation of the hierarchical open-vocabulary scene graphs (Sec. III-B), and (iii) language-conditioned navigation across large-scale environments (Sec. III-C). Fig. 2 presents an overview of our method.
@@
III-A1 Frame-Wise 3D Segment Merging
Given a sequence of RGB-D observations, we utilize Segment Anything [21] to obtain a list of class-agnostic 2D binary masks at each timestep. The pixels in each mask are then backprojected to 3D using the depth information, resulting in a list of point clouds, or 3D segments. Based on accurate odometry estimates, we transform all 3D segments into the global coordinate frame. These frame-wise segments are either initialized as new global segments or merged with existing ones based on an overlap metric:
where and indicate segment (or point cloud) and , is computed by taking the number of points in showing a neighbor in within a certain distance divided by the total number of points in . Different from Gu et al. [11], who incrementally merge new segments with one global segment that has the largest overlapping ratio, we construct a graph with edge weights according to the corresponding overlapping ratios. Based on these weights, highly-connected subgraphs are subsequently merged. In this way, one segment can merge with more segments, which is useful in situations in which an incoming segment is, e.g., filling a gap between two already registered global segments.
III-A2 Open-Vocabulary Segment Features
For each obtained 2D SAM mask per frame, we obtain an image crop based on its bounding box as well as an image of the isolated mask without background. We encode the full RGB frame and the two mask-wise images with CLIP [20] and fuse them in a weighted-sum manner, see Fig. 2. Previous work [39] proposed to use the CLIP feature of the masked image without background, while others [10] approach this by combining the CLIP features of the whole image and the target mask’s crop including background. In our work, we empirically show that encoding the full RGB frame and the two mask-wise images with CLIP and fusing them in a weighted-sum manner achieves improved results. The fusion scheme can be formulated as:
@@ III-A2 Open-Vocabulary Segment Features
Assuming a single CLIP feature for each mask, we transform the 2D mask into global 3D coordinates and associate the obtained fused CLIP feature with the nearest 3D points in a pre-computed reference point cloud. Based on this association, we register the obtained segment features on a global point-wise feature map. The final point-wise features are then determined by averaging each reference point’s associated features. Based on the 3D segments obtained in the independent merging step, we can finally infer open-vocabulary vision-language features for all 3D segments as outlined in Fig.2. In the subsequent step, we match point-wise features with the obtained 3D segments. For each point within a segment, we identify the nearest points in the reference point cloud and collect their CLIP features. Thus, we obtain a set of various CLIP features per 3D segment that is turned into a single segment-wise feature as follows: To mitigate potential high variance, instead of directly averaging these features, we employ DBSCAN clustering to handle cases like under-segmentation or imperfect views, ensuring a more representative feature by selecting the feature closest to the mean of the majority cluster.
@@
In this section, we describe how to build a hierarchical open-vocabulary scene graph given a global reference point cloud of the scene, a list of global 3D segments, and their associated CLIP features as described in Sec. III-A. We formalize our graph as where denotes the nodes and denotes the edges. The nodes can be expressed as , consisting of a root node , floor nodes , room nodes , and object nodes . Each node in the graph except the root node contains the point cloud of the concept it refers to and the open-vocabulary features associated with it. The edges can be written as . Here, represents the edges between the root node and the floor nodes, represents the edges between the floor nodes and the room nodes, and lastly, denotes the edges between the room and object nodes.
III-B1 Floor Segmentation
@@
In order to separate floors, we identify peaks within a height histogram over all points contained in the point cloud. Given the point cloud of the whole environment, we construct the histogram over all points along the height axis using a bin size . Next, we identify peaks in this histogram (within a local range of ) and select only peaks that exceed a minimum of 90% of the highest intensity peak. We apply DBSCAN and select the two highest-ranking peaks in each cluster. After that, every two consecutive values in the sorted height vector represent a single floor (floor and ceiling) in the building. The floor segmentation process is shown in Fig. 3. Using the obtained height levels, we can extract floor point clouds for each floor where is the floor number. In addition, we equip each floor node with a CLIP text embedding using the template “floor {#}”. A graph edge between the root node and this floor node is also established.
III-B2 Room Segmentation
@@
Based on each obtained floor point cloud, we construct a 2D bird’s-eye-view (BEV) histogram, from which the binary wall skeleton mask is extracted by thresholding the histogram. After dilating the wall mask and computing an Euclidean distance field (EDF), a number of isolated regions is derived by thresholding the EDF. Taking these regions as seeds, we apply the Watershed algorithm to obtain 2D region masks. The room segmentation process is further shown in Fig. 3. Given the 2D region masks, we extract the 3D points that fall into the floor’s height interval as well as the BEV room segment to form room point clouds that are used to associate objects to rooms later.
@@
In order to attribute room nodes, we associate RGB-D observations whose camera poses reside within a room segment to those rooms, see Fig. 4. The CLIP embeddings of these images are filtered by extracting representative view embeddings using the k-means algorithm. Given a list of room categories encoded via CLIP, we construct a cosine similarity matrix between the representative features and all room category features. Next, we take the along the category axis and obtain the most probable room type for each representative separately, resulting in votes per room. Given these votes, we obtain the predicted room category by either taking the maximum-score vote or the majority vote across all representatives per room.
@@
These representative embeddings and the room point cloud attribute the room node , representing room on floor . An edge between the floor node and each room node is established. The construction and querying of view embeddings are illustrated in Fig. 4.
III-B3 Object Identification
@@
Given the room point cloud, we associate object-level 3D segments that show a point cloud overlap with a potential candidate room in the bird’s-eye-view. Whenever a segment shows zero overlap with any room, we associate it with the room showing the smallest Euclidean distance. To reduce the number of nodes, we merge 3D segments of significant pair-wise partial overlap (Sec. III-A1) that produce equal object labels when queried against a chosen label set. Each merged point cloud constitutes an object node that is connected to its corresponding room node by an edge . Each object node holds its corresponding 3D segment feature as described in Sec. III-A1, its 3D segment point cloud as well as a maximum-score object label for intermediate naming.
III-B4 Actionable Navigational Graph
@@
In addition to the open-vocabulary hierarchy, the scene graph also contains a navigational Voronoi graph that serves robotic traversability of the mapped surroundings [40] spanning multiple floors. This enables high-level planning and low-level execution based on the Voronoi graph. The creation of actionable graphs involves constructing per-floor and cross-floor navigation graphs. For the floor-level graph, the approach entails computing the free space map of the floor and creating a Voronoi graph [40] based on it. To construct per-floor graphs, we first obtain all camera poses and project them as 2D points onto a BEV map of each floor, assuming areas within a certain radius of each pose are pair-wise navigable. Subsequently, the entire floor’s region is obtained by projecting all floor-wise into the BEV plane. An obstacle map is generated based on points within a predefined height range , where is the minimal height of the floor points and , are two thresholds we identify manually as and . By taking the union of the pose region map and the floor region map and subtracting the obstacle region map, the free space map of each floor is derived. The Voronoi graph of this free map yields the floor graph. (See Fig. 5(a)). To build cross-floor navigation graphs, camera poses on stairs are connected to form a stairs-wise graph. Then, the closest nodes between the stairs graph and the floor-wise graph are selected and connected, completing the construction of cross-floor navigational graphs as shown in Fig. 5(b).
III-C Object Navigation from Long Queries

### v2

@@ Abstract
Recent open-vocabulary robot mapping methods enrich dense geometric maps with pre-trained visual-language features. While these maps allow for the prediction of point-wise saliency maps when queried for a certain language concept, large-scale environments and abstract queries beyond the object level still pose a considerable hurdle, ultimately limiting language-grounded robotic navigation. In this work, we present HOV-SG, a hierarchical open-vocabulary 3D scene graph mapping approach for language-grounded indoor robot navigation. Leveraging open-vocabulary vision foundation models, we first obtain state-of-the-art open-vocabulary segment-level maps in 3D and subsequently construct a 3D scene graph hierarchy consisting of floor, room, and object concepts, each enriched with open-vocabulary features. Our approach is able to represent multi-story buildings and allows robotic traversal of those using a cross-floor Voronoi graph. HOV-SG is evaluated on three distinct datasets and surpasses previous baselines in open-vocabulary semantic accuracy on the object, room, and floor level while producing a 75% reduction in representation size compared to dense open-vocabulary maps. In order to prove the efficacy and generalization capabilities of HOV-SG, we showcase successful long-horizon language-conditioned robot navigation within real-world multi-story environments. We provide code and trial video data at: https://hovsg.github.io.
@@ I Introduction
Humans acquire conceptual knowledge about the world as a whole and concrete objects in particular through multi-modal experiences. These semantic experiences are paramount to object recognition and language as well as reasoning and planning [1, 2]. Cognitive maps store this information based on sensor fusion, fragmentation, and hierarchical structure. This is central to the human ability to navigate the physical world [3, 4, 5]. Recently, language proved to be an effective link between intelligent systems and humans and can enable robot autonomy in complex human-centered environments [6, 7, 8, 9, 10, 11, 12, 13, 14].
Classical methods for robot navigation build dense spatial maps of high accuracy using approaches to simultaneous localization and mapping (SLAM) [15]. Those give rise to fine-grained navigation and manipulation based on geometric goal specifications. Recent advances have combined dense maps with pre-trained zero-shot vision-language models, which facilitates open-vocabulary indexing of observed environments [9, 10, 11, 16, 13, 17, 12]. While these approaches marry the area of classical robotics with modern open-vocabulary semantics, representing larger scenes while abstracting still poses a considerable hurdle. Scalable scene representations generated from real-world perception inputs should generally fulfill the following requirements: 1) Object-centricity and abstraction in terms of hierarchies, 2) efficiency regarding storage capacity as well as actionability, 3) true open-vocabulary indexing and easy querying.
@@
In this work, we present Hierarchical Open-Vocabulary 3D Scene Graphs, short HOV-SG. Our approach abstracts from dense open-vocabulary maps and allows the indexing of three distinct concepts, namely floors, rooms as well as objects. We utilize open-vocabulary vision-language models [21, 22] across all concepts in order to construct 3D scene graph hierarchies that span multi-story environments while maintaining a small memory footprint. Given its concept-centric nature, the HOV-SG representation is promptable using LLMs. Different from previous work, our approach relates to different conceptual levels by first decomposing abstract queries such as “towel in the bathroom on the upper floor” and scoring the obtained tokens against the different levels of the hierarchy. We complement this with a navigational Voronoi graph that covers multiple floors including stairs, which allows actionable grounding of decomposed queries in the environment. This enables object retrieval and long-horizon robotic navigation in large-scale indoor environments from abstract queries as shown in Fig. 1.
In summary, we make the following contributions:
1.
We introduce a novel fusion scheme using feature clustering of zero-shot embeddings that yields state-of-the-art results in open-set 3D semantic segmentation.
2.
We present an algorithm that enables the construction of truly actionable open-vocabulary 3D scene graphs of multi-floor buildings.
- 3.
4.
We conduct real-world multi-floor object navigation experiments based on long natural language queries.
5.
We introduce a novel evaluation metric for measuring open-vocabulary semantics termed AUC.
6.
We make our code and evaluation protocol publicly available at https://hovsg.github.io to foster future research and introduce comparability in open-set mapping.
@@ II-A Semantic 3D Mapping
3D scene graphs have emerged as an effective, object-centric representation of large-scale indoor [36, 19, 20] and outdoor scenes [18]. By representing objects or spatial concepts as nodes and their relations as edges, 3D scene graphs allow to efficiently represent larger scenes [36, 18, 19]. Both edges and nodes can hold geometric and semantic attributes, which are often inferred from certain off-the-shelf networks [37]. Decomposing scenes into objects and their relations enables higher-level reasoning for robotic navigation and manipulation. This is particularly useful in the realm of reasoning, planning, and navigation given the object-centric nature of these tasks [19, 38]. Often combined with odometry estimates from simultaneous localization and mapping (SLAM) [39, 18, 40], 3D scene graphs also allow a tight coupling between semantics and highly accurate mapping approaches utilizing e.g. meshes to represent environments [19].
@@
Early works have shown how to encode hierarchies via abstraction in both the spatial and the semantic domain using offline approaches [36, 20]. Successive works investigated learning-based scene graph construction [41, 37] as well as dynamic indoor scenes [19]. Several approaches such as SceneGraphFusion [37] and S-Graphs [39] also investigate the real-time capabilities of their proposed approaches. Most recently, ConceptGraphs [14] was the first to show how to combine 3D scene graphs with open-vocabulary vision-language features. In addition, the authors show how to query the graph using LLMs and demonstrate various downstream applications.
II-C Scene Graphs for Planning
Several recent works have investigated the use of scene graphs for robotic planning. The earliest approaches rely on pre-explored environments and perform iterative scene graph decomposition to retrieve grounded plans [38, 42]. RobLM [42] decomposes the planning stage by relying on a fine-tuned GPT-2 instance that proposes high-level sub-problems from scene graphs, which are in turn solved through PDDL task planners. SayPlan [38] directly utilizes GPT-4 [43] for iterative search on a scene graph to generate grounded plans, which requires feasibility constraints on the manipulated entities and actions. Another line of work investigates robotic navigation from scene graphs. SayNav [44] obtains LLM-generated plans from scene graphs and executes short-distance point-goal navigation sub-tasks. Contrary to that, VoroNav [45] constructs a Voronoi graph that is attributed to camera observations in order to solve object navigation. Orthogonal to that, MoMa-LLM [46] tackles mobile manipulation objectives using scene graphs fed to GPT-4 in a task-specific manner. Similarly, GRID [47] uses a graph neural network to predict actions from scene graphs and LLM encodings.
@@ II-C Scene Graphs for Planning
Conceptually, our work is most similar to ConceptGraphs and Hydra. While ConceptGraphs is only evaluated on small scenes and mostly validated by human evaluators in terms of semantic accuracy of nodes etc., our work proposes not only a novel metric for measuring the semantic accuracy of object features but also introduces open-vocabulary hierarchies. Different from ConceptGraphs and similar to Hydra [19], which does not operate on open-vocabulary features, our approach demonstrates how to efficiently represent actionable, hierarchical 3D scene graphs that are attributed with open-vocabulary features.
@@ III Technical Approach
This work aims to develop a concise and efficient visual-language graph representation for large-scale multi-floor indoor environments given RGB-D observations and odometry. The graph should facilitate the indexing of multi-level semantic concepts through natural language queries such as “the first floor” (floor level), “the office on the first floor” (room level), and “the plant in the office on the second floor” (object level). Additionally, the graph should be actionable and enable a robot to localize and navigate semantically and spatially in the environment without additional geometric maps. We address this by introducing Hierarchical Open-Vocabulary Scene Graphs, in short HOV-SG. The overall pipeline consists of two stages. We first create a 3D segment-level open-vocabulary map and then build a hierarchical open-vocabulary scene graph based on the map. In the following sections, we describe (i) the construction of the 3D segment-level open-vocabulary map (Sec. III-A), (ii) the creation of the hierarchical open-vocabulary scene graph (Sec. III-B), and (iii) how to use the graph for language-conditioned navigation across a large-scale environment (Sec. III-C). Fig. 2 presents an overview of our method.
@@
The main idea of building a segment-level open-vocabulary map is to create a list of 3D point clouds, namely segments, from an RGB-D video with odometry and assign an open-vocabulary feature generated by a pre-trained visual-and-language model (VLM) to each segment. Unlike previous works that equip each 3D point with an independent visual-language feature [11, 10, 13, 12], we leverage the fact that neighboring points in the 3D world often share the same semantic information. This implies the potential of reducing the required semantic features to represent the scene while maintaining expressiveness.
Frame-Wise 3D Segment Merging: Given a sequence of RGB-D observations, we utilize Segment Anything [22] to obtain a list of class-agnostic 2D binary masks at each timestep. The pixels in each mask are then backprojected to 3D using the depth information, resulting in a list of point clouds, or 3D segments. Based on accurate odometry estimates, we transform all 3D segments into the global coordinate frame. These frame-wise segments are either initialized as new global segments or merged with existing ones based on an overlap metric:
where and indicate segment (or point cloud) and , is computed by taking the number of points in showing a neighbor in within a certain distance divided by the total number of points in . Different from Gu et al. [14], who incrementally merge new segments with one global segment that has the largest overlapping ratio, we construct a fully connected graph where each segment serves as a node and their edge weights are the corresponding overlapping ratios. Based on these weights, highly-connected subgraphs are subsequently merged. In this way, one segment can be merged with multiple segments, which is useful in situations in which an incoming segment is, e.g., filling a gap between two already registered global segments.
Segment-Level Open-Vocabulary Features Computation: For each obtained 2D SAM mask per frame, we obtain an image crop based on its bounding box as well as an image of the isolated mask without background. We encode the full RGB frame and the two mask-wise images with CLIP [21] and fuse them in a weighted-sum manner (Fig. 2, left). Previous work [48] proposed to use the CLIP feature of the masked image without background, while others [13] approach this by combining the CLIP features of the whole image and the target mask’s crop including background. In our work, we empirically show that encoding the full RGB frame and the two mask-wise images with CLIP and fusing them in a weighted-sum manner achieves improved results (Sec. IV-E). The fusion scheme can be formulated as:
@@
Assuming a single CLIP feature for each mask, we transform the 2D mask into global 3D coordinates and associate the obtained fused CLIP feature with the nearest 3D points in a pre-computed reference point cloud. Based on this association, we register the obtained segment features on a global point-wise feature map. The final point-wise features are determined by averaging each reference point’s associated features. Based on the 3D segments obtained in the independent merging step, we can finally infer open-vocabulary vision-language features for all 3D segments as outlined in Fig.2. In the subsequent step, we match point-wise features with the obtained 3D segments. For each point within a segment, we identify the nearest points in the reference point cloud and collect their CLIP features. We leverage DBSCAN to cluster all the point-wise features of the segment and assign the feature that is closest to the majority cluster’s mean to the segment (Fig. 2, middle). This circumvents mode collapse while removing noise and thus produces more semantically meaningful segment features.
@@
In this section, we describe how to build a hierarchical open-vocabulary scene graph given a global reference point cloud of the scene, a list of global 3D segments, and their associated CLIP features as described in Sec. III-A.
We formalize our graph as where denotes the nodes and denotes the edges. The nodes can be expressed as , consisting of a root node , floor nodes , room nodes , and object nodes . Each node in the graph except the root node contains the point cloud of the concept it refers to and the open-vocabulary features associated with it. The edges can be written as . Here, represents the edges between the root node and the floor nodes, represents the edges between the floor nodes and the room nodes, and lastly, denotes the edges between the room and object nodes.
@@
Floor Segmentation: In order to separate floors, we identify peaks of a height histogram over all points contained in the point cloud. Given the point cloud of the whole environment, we construct the histogram over all points along the height axis using a bin size . Next, we identify peaks in this histogram (within a local range of ) and select only peaks that exceed a minimum of 90% of the highest intensity peak. We apply DBSCAN and select the two highest-ranking peaks in each cluster. After that, every two consecutive values in the sorted height vector represent a single floor (floor and ceiling) in the building. The floor segmentation process is shown in Fig. 3. Using the obtained height levels, we can extract floor point clouds for each floor where is the floor number. In addition, we equip each floor node with a CLIP text embedding using the template “floor {#}”. A graph edge between the root node and each floor node is established.
@@
Room Segmentation: Based on each obtained floor point cloud, we construct a 2D bird’s-eye-view (BEV) histogram, from which a binary wall skeleton mask is extracted by thresholding the histogram. After dilating the wall mask and computing an Euclidean distance field (EDF), a number of isolated regions is derived by thresholding the EDF. Taking these regions as seeds, we apply the Watershed algorithm to obtain 2D region masks. The room segmentation process is further shown in Fig. 3. Given the 2D region masks, we extract the 3D points that fall into the floor’s height interval as well as the BEV room segment to form room point clouds that are used to associate objects to rooms later.
@@
To enrich room nodes with open-vocabulary features, we associate RGB-D observations whose camera poses reside within a room segment to those rooms (see Fig. 4). The CLIP embeddings of these images are distilled by extracting representative view embeddings using the k-means algorithm. During inference, given a list of room categories encoded via CLIP, we construct a cosine similarity matrix between the representative features and all room category features. Next, we take the along the category axis and obtain the most probable room type for each representative separately, resulting in votes per room. Given these votes, we obtain the predicted room category by either taking the maximum-score vote or the majority vote across all representatives per room.
@@
These representative embeddings and the room point cloud attribute the room node of room on floor . An edge between the floor node and each room node is established. The construction and querying of view embeddings are illustrated in Fig. 4.
@@
Object Identification: Given the room point cloud, we associate object-level 3D segments that show a point cloud overlap with a potential candidate room in the bird’s-eye-view. Whenever a segment shows zero overlap with any room, we associate it with the room showing the smallest Euclidean distance. To reduce the number of nodes, we merge 3D segments of significant pair-wise partial overlap (Sec. 3) that produce equal object labels when queried against a chosen label set. Each merged point cloud constitutes an object node that is connected to its corresponding room node by an edge . Each object node holds its corresponding 3D segment feature as described in Sec. III-A, its 3D segment point cloud as well as a maximum-score object label for intermediate naming.
@@
Actionable Graph Creation: In addition to the open-vocabulary hierarchy, the scene graph also contains a navigational Voronoi graph that serves robotic traversability of the mapped surroundings [49] spanning multiple floors. This enables high-level planning and low-level execution based on the Voronoi graph. The creation of actionable graphs involves constructing per-floor and cross-floor navigation graphs. For the floor-level graph, the approach entails computing the free space map of the floor and creating a Voronoi graph [49] based on it. To construct per-floor graphs, we first obtain all camera poses and project them as 2D points onto a BEV map of each floor, assuming areas within a certain radius of two nodes are pair-wise navigable. Subsequently, the entire floor’s region is obtained by projecting all floor-wise points into the BEV plane. An obstacle map is generated based on points within a predefined height range , where is the minimal height of the floor points while , are empirically tuned. By taking the union of the pose region map and the floor region map and subtracting the obstacle region map, the free space map of each floor is derived. The Voronoi graph of this free map yields the floor graph. (see Fig. 5(a)). To build cross-floor navigational graphs, camera poses on stairs are connected to form stair-wise graphs. Subsequently, the closest nodes between the stairs graph and the floor-wise graph are selected respectively and connected, thereby completing the construction of cross-floor navigational graphs as shown in Fig. 5(b).
III-C Navigation with Scene Graph